In an inelastic collision, two objects collide with each other and come to a rest. The first object’s initial momentum was 30kg⋅m/s . If the second object’s mass is 40kg , what was its initial momentum?

a) −30kg⋅m/s

b) 0kg⋅m/s

c) 30kg⋅m/s

d) 40kg⋅m/s
Answer #1

Since in inelastic collision too the momentum is conserved,

so, m1 x v1 + m2 x v2 = (m1+m2) x 0   (as after collision both objects come to rest)

m1 x v1 = 30 kg.m/s (given)

m2 = 40 kg (given)

v2 = -30/40

so its initial momentum = m2 x v2 = 40 x (-30/40) = -30 kg.m/s